# Temporal-Spectral-Spatial Unified Remote Sensing Dense Prediction

### v1

@@ VI CONCLUSION
This paper addressed the critical challenge of processing heterogeneous remote sensing data for diverse dense prediction tasks, a common hurdle due to variations in temporal, spectral, and spatial characteristics. We introduced the Temporal-Spectral-Spatial Unified Network, a novel deep learning architecture. TSSUN demonstrates remarkable flexibility by handling arbitrary input configurations across these three dimensions and unifying key dense prediction tasks: semantic segmentation, semantic change detection, and binary change detection, with adaptable output class settings. The proposed Temporal-Spectral-Spatial Unified Strategy enables this unification by effectively decoupling and modeling data representations guided by meta-information. Furthermore, the designed Local-Global Window Attention mechanism significantly improves feature extraction by capturing both local and global contextual information efficiently. Comprehensive experimental validation on multiple datasets confirmed that a single TSSUN model consistently adapts to varied data inputs and task requirements, achieving or exceeding state-of-the-art performance. The TSSUN framework represents a substantial advancement towards universal remote sensing models, mitigating the need for task-specific designs and extensive retraining, thus paving the way for more scalable and cost-effective analysis of complex, multi-source geospatial data and fostering broader application of remote sensing technologies.

### v2

@@ VI CONCLUSION
This study introduced the Spatial-Temporal-Spectral Unified Network, a novel architecture designed to overcome the critical limitations of model rigidity in remote sensing dense prediction. By establishing a unified representation for arbitrary spatial, temporal, and spectral data configurations, STSUN demonstrates exceptional adaptability to heterogeneous datasets. STSUN uniquely integrates semantic segmentation, binary change detection, and semantic change detection within a single model via trainable task embeddings, obviating the need for specialized, task-specific architectures. Moreover, STSUN utilizes trainable semantic category embeddings to perform dense prediction for any prediction class setting without requiring model retraining. Comprehensive experimental validation on multiple datasets confirmed that a single STSUN model consistently adapts to varied data inputs and outputs, unifing multiple dense prediction tasks and prediction category settings, achieving or exceeding state-of-the-art performance. We envision STSUN to serve a baseline for universal remote sensing dense prediction models, mitigating the need for task-specific designs and extensive retraining.